OpenStudy (anonymous):

Solve the system of equations given below. 5x + 4y = 25 3x + 4y = 15

OpenStudy (anonymous):

Plz help ps below of the the answers (5, 0) (0, 0) (0, -5) (-5, -5)

OpenStudy (katrinakaif):

Distribute a negative to the second equation since then you can cancel out +4y and -4y = 0y So: 5x + 4y = 25 -(3x + 4y) = (15)- = -3x - 4y = -15 Now simplify it : 5x - 3x = 2x + 4y - 4y = 0y 25 - 15 = 10 You have an equation of : 2x = 10 x = 5 Plug the x into one of the equations to get your "Y" =)

OpenStudy (anonymous):

y = 0 :D

OpenStudy (anonymous):

(5, 0)